A turboprop engine consists of a diffuser, compressor, combustor, turbine, and nozzle. The turbine drives a propeller as well as the compressor. Air enters the diffuser with a volumetric flow rate of 63.7 m3/s at 40 kPa, 240 K, and a velocity of 180 m/s, and decelerates essentially to zero velocity. The compressor pressure ratio is 9 and the compressor has an isentropic efficiency of 85%. The turbine inlet temperature is 1240 K, and its isentropic efficiency is 85%. The turbine exit pressure is 50 kPa and the nozzle exit pressure is 40 kPa. Flow through the diffuser and nozzle is isentropic. Use an air-standard analysis and neglect kinetic energy except at the diffuser inlet and the nozzle exit.Determine the power delivered to the propeller, in MW.
Answer #1

W, = 9.2MW

The detailed solution is given below:
Giveno va = 63.7 mils | Ta=200k [v=o | Me = M = 0.85 Pa= 40kba / Va = 180m/s - 9 / T = 1240K Compressor Combustor o 1 Pq = 50

0. At To =25ok → ho = 250.05 and Port = 0.7329 and TC= 260k → hc = 260.09 and Prc = 0.8405 → Pro = Pro + (re-Pro) x[h-ho] = 0

c. b = 2502 + [ato ] seos king # State 3: [P3=] - - To=1240k h ₂ = 1324.93 kilky I paz = 272.3 # Stalen: [isen tropic expansi

= 1324.93 – (0.85) (1324-93–722) 6. h4 = 812.44 kilky # The power delivered to the propellers → Wp = WE Wc = in [[hz ha) -(hz